If ex-radii r1, r2, r3 of a triangle ABC are in H.P., then the sides of the triangle are in
A. A.P.
B. G.P.
C. H.P.
D. none of the above

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Using the relations
r1 = S/(s-a) where S = area and s is semi-perimeter.
r2 = S/(s-b)
r3 = S/(s-c).
As the radii are in HP their reciprocals are in A.P.
So (s-a)/S,..,.. are in AP or s-a, s-b, s-c are in AP so a,b,c are in AP
Hence A is the correct answer
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty